Sampson Corp. is evaluating the introduction of a new product. The possible levels of unit sales and the probabilities of their occurrence are shown next:
  

Possible
Market Reaction

Sales
in Units
Probabilities
Low response 70 .40
Moderate response 100 .30
High response 120 .10
Very high response 150 .20


a. What is the expected value of unit sales for the new product? (Do not round intermediate calculations and round your answer to the nearest whole unit.)
  


  
b. What is the standard deviation of unit sales? (Do not round intermediate calculations. Round your answer to 2 decimal places.)
Answer #1

Expected Sales=Respective Sales*Respective probability

=(0.4*70)+(0.3*100)+(0.1*120)+(0.2*150)=100 units

probability Sales probability*(Sales-Expected Sales)^2
0.4 70 0.4*(70-100)^2=360
0.3 100 0.3*(100-100)^2=0
0.1 120 0.1*(120-100)^2=40
0.2 150 0.2*(150-100)^2=500
Total=900

Standard deviation=[Total probability*(Sales-Expected Sales)^2/Total probability]^(1/2)

=30 units.